OpenStudy (anonymous):

derivative question. (a) find an equation of the tangent line to the graph of y=g(x) at x=5 if g(5)=-3 and g'(5)=4 (b)if the tangent line to y=f(x) at (4,3) passes through the point (0,2), find f(4) and f'(4)

OpenStudy (anonymous):

y + 3 = 4 ( x -5) => y = 4x - 23 f ( 4) = 3 -> f'( 4) = 0 y - 2 = f'(4) x

OpenStudy (rogue):

Part A is correct. f(x) includes the point (4,3), so f(4) = 3. f'(4) is the slope of the tangent, which is (3-2)/(4-0) = 1/4.

OpenStudy (anonymous):

I'm thinking of how to explain, but it's taking me forever, then I finally give it up!